OpenStudy (anonymous):

Myininaya's great new method for factoring quadratics (please don't award medals in this thread, u can award them in the below thread if u want to)

OpenStudy (anonymous):

http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e3a5e000b8bf47d066015a8 here are some examples here ax2+bx+c you want to find two factors of a*c that have product a*c and have sum b so we have b=(b/2+z)+(b/2-z) a*c=(b/2+z)(b/2-z)=b^2/4-z^2 <- solve this for z then replace bx with (b/2+z)x+(b/2-z)x then factor by grouping :)

hero (hero):

Show an example using a real quadratic....like x^2 + 7x-1 or something

hero (hero):

:o)

myininaya (myininaya):

ok b=7=(7/2+z)+(7/2-z) a*c=-1=(7/2+z)(7/2-z) -1=49/4-z^2 z^2=49/4+1 z^2=53/4 so \[z=\sqrt{\frac{53}{4}}\] so we have \[bx=(7/2+\frac{\sqrt{53}}{2})x+(7/2-\frac{\sqrt{53}}{2})x\] so we have \[x^2+(7/2+\frac{\sqrt{53}}{2})x+(7/2-\frac{\sqrt{53}}{2})x-1\] now factor by grouping \[x(x+\frac{7+\sqrt{53}}{2})+\frac{7-\sqrt{53}}{2}(x-\frac{1*2}{7-\sqrt{53}})\] \[x(x+\frac{7+\sqrt{53}}{2})+\frac{7-\sqrt{53}}{2}(x-\frac{2(7+\sqrt{53})}{49-53})\] \[x(x+\frac{7+\sqrt{53}}{2})+\frac{7-\sqrt{53}}{2}(x+\frac{7+\sqrt{53}}{2})\] \[(x+\frac{7+\sqrt{53}}{2})(x+\frac{7-\sqrt{53}}{2})\]

myininaya (myininaya):

:)

hero (hero):

Thanks much :)

OpenStudy (anonymous):

Beautiful Solution

myininaya (myininaya):

its pretty long but i love it :)

hero (hero):

It's awesome....

hero (hero):

Why did you use z instead of d

hero (hero):

I'm nitpicking

myininaya (myininaya):

you can use d lol i just got stuck on z because satellite said z

myininaya (myininaya):

i was using ? instead

hero (hero):

Oh

myininaya (myininaya):

z is a nicer looking variable then ?

OpenStudy (anonymous):

Ok, fess up...who awarded the medal?

hero (hero):

Me? How many cool points do I lose for it?

hero (hero):

Myininaya, you just proved that any quadratic can be factored.

myininaya (myininaya):

yep i did. it also works for the complex number :)

myininaya (myininaya):

i mean i can factor over complex numbers too :)

hero (hero):

Excellent. I'm having an argument with a tutor over it as we speak. You should publish this somewhere....

myininaya (myininaya):

an example over the complex numbers

myininaya (myininaya):

how do i publish this

hero (hero):

Magnificent....

hero (hero):

I found this: www.ams.org/notices/200711/tx071101507p.pdf

hero (hero):

Honestly, I feel like I've been lied to my whole life....

myininaya (myininaya):

lol why

myininaya (myininaya):

because you didn't know you could factored all quadratics without the quadratic formula

OpenStudy (anonymous):

used z because i used it!!

myininaya (myininaya):

thats what i said lol

OpenStudy (anonymous):

lol

myininaya (myininaya):

i gave you credit

myininaya (myininaya):

i could chose some other letter or stuck with ?

OpenStudy (anonymous):

i don't deserve it. make a nice pdf of this and save it

myininaya (myininaya):

but i wanted to please you

OpenStudy (anonymous):

thank you

myininaya (myininaya):

ok of the general thing right?

myininaya (myininaya):

satellite i came up with something i have a brain omg this is so weird

OpenStudy (amistre64):

nice job :) but its just not abstract enough for me ;)

OpenStudy (amistre64):

needs more celtic runes, and a dash of thyme

myininaya (myininaya):

lol

myininaya (myininaya):

http://www.purplemath.com/modules/quadform.htm this site sometimes the quadratic doesn't factor at all lol

myininaya (myininaya):

myininaya>purplemath

hero (hero):

What the.....lol

hero (hero):

What does Purple Math mean anyway?

myininaya (myininaya):

i dont know what that means lol i'm trying to see if i can find what i have done somewhere

hero (hero):

Probably not

myininaya (myininaya):

here's another example if anyone is interested: \[x^2+1\] \[x^2+0x+1\] a=1,b=0,c=1 \[b=(\frac{0}{2}-z)+(\frac{0}{2}+z)\] \[a*c=1=(\frac{0}{2}-z)(\frac{0}{2}+z)=-z^2\] \[z=i\] \[x^2+-ix+ix+1\] \[x(x-i)+i(x+\frac{1}{i})=x(x-i)+i(x+\frac{1}{i}*\frac{i}{i})\] \[x(x-i)+i(x+\frac{i}{-1})\] \[x(x-i)+i(x-i)=(x-i)(x+i)\]

hero (hero):

Much simpler

myininaya (myininaya):

well we could had solved x^2=-1 to begin with lol

hero (hero):

hmmm

myininaya (myininaya):

to find the factors

hero (hero):

true x^2 = -1

myininaya (myininaya):

x=i , x=-i x-i=0, x+i=0 which means x^2+1=(x-i)(x+i)

hero (hero):

beautiful

OpenStudy (zarkon):

very nice!

myininaya (myininaya):

it is very cool right? i couldn't find anything like it on the internet yet is it unique i wonder?

OpenStudy (zarkon):

I don't know

myininaya (myininaya):

so we don't have to use completing the square to prove the quadratic formula

OpenStudy (zarkon):

it gives another nice way to derive the quadratic equation :)

OpenStudy (zarkon):

good stuff myininaya

myininaya (myininaya):

i made a mistake in my pdf

myininaya (myininaya):

i think

OpenStudy (anonymous):

try latex much nicer

myininaya (myininaya):

dang it i see the mistake

myininaya (myininaya):

i wrote a stupid a where there shouldn't had been one i been really lame at math lately

myininaya (myininaya):

i was so excited i sent all of this to my ex-teachers and i didn't make it pretty or check it lol

OpenStudy (zarkon):

I've done that before with students :0

myininaya (myininaya):

lol

myininaya (myininaya):

i hope they will be interested in my emails

OpenStudy (zarkon):

I'm sure they will

myininaya (myininaya):

:)

OpenStudy (anonymous):

you want a latex version?

OpenStudy (zarkon):

now you need to generalize your technique to cubic polynomials ;)

myininaya (myininaya):

ha ha

hero (hero):

Yes, I was thinking that too...and perhaps even for all polynomials

myininaya (myininaya):

lol

myininaya (myininaya):

you guys are crazy

myininaya (myininaya):

especially you hero

OpenStudy (zarkon):

that would not be possible :)

hero (hero):

At least for cubic polynomials

myininaya (myininaya):

i think they might have a way to solve cubics but it isn't popular for some reason and i don't know about it

OpenStudy (zarkon):

cubics and quartics, but that is it

myininaya (myininaya):

why don't they talk about those in algebra

myininaya (myininaya):

i mean what do you need to solve them? it is just algebra?

myininaya (myininaya):

omg

OpenStudy (anonymous):

@ zarkon i found a formula for 5th degree polynomial. it is very elegant

OpenStudy (zarkon):

really?

myininaya (myininaya):

i dont trust that word elegant when the cubic wasn't elegant

OpenStudy (anonymous):

that was my feeble attempt at a joke. galois says it is not possible.

myininaya (myininaya):

lol

OpenStudy (zarkon):

"However, there is no formula for general quintic equations over the rationals in terms of radicals; this is known as the Abel–Ruffini theorem, first published in 1824,"

OpenStudy (zarkon):

;)

OpenStudy (zarkon):

A5 is simple ;)